### v2

# Passing the Buck to AI: How Individuals’ Decision-Making Patterns Affect Reliance on AI

Abstract.
Psychological research has identified different patterns individuals have while making decisions, such as vigilance (making decisions after thorough information gathering), hypervigilance (rushed and anxious decision-making), and buckpassing (deferring decisions to others). We examine whether these decision-making patterns affect peoples’ engagement with AI-generated information in decision-making. In an online experiment with 810 participants tasked with distinguishing food facts from myths, we found that a higher buckpassing tendency was positively correlated with the likelihood of seeking AI information and reported reliance on AI, while being negatively correlated with the time spent reading AI explanations. In contrast, the higher a participant tended towards vigilance, the more carefully they scrutinized the AI’s information, as indicated by an increased time spent looking through the AI’s explanations. These findings suggest that a person’s decision-making pattern plays a significant role in their interactions with AI suggestions, which provides a new understanding of individual differences in AI-assisted decision-making.
decision-making, user characteristics, psychology, Human-AI interaction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Conversational AI tools like ChatGPT provide new ways for individuals to obtain information that can support their decision-making. With their widespread availability and advanced capabilities, people increasingly use these tools to ask for nutrient suggestions (Niszczota and Rybicka, 2023), medical advice (Liu et al., 2023), or personal financial recommendations (Hassani and Silva, 2023). Despite their extensive capabilities, current conversational AI tools supported by large language models (LLMs) have been found to generate inaccurate and biased information in various domains including politics, law, and medicine (Bubeck et al., 2023; Dahl et al., 2024; Guerreiro et al., 2023; Tao et al., 2023; Rozado, 2023; Niszczota and Rybicka, 2023). Reliance on these tools without caution could therefore lead to misinformed decisions and contribute to the spread of misinformation (Feng et al., 2023).
To mitigate these risks and inform interventions, research in Human-AI interaction (HAI) has focused on factors related to AI systems that affect users’ reliance, such as accuracy and transparency (Kim et al., 2023; Cao et al., 2024; Zhang et al., 2020; Wang and Yin, 2021; Schemmer et al., 2023; Schoeffer et al., 2024; Wang and Yin, 2023). Other research has explored how individual differences are correlated with varying behaviors in HAI. For example, researchers have shown that attitudes toward, and reliance on, AI vary across different demographics and personality traits (Lai and Tan, 2019; Buçinca et al., 2021; Kaya et al., 2024; Ma and Huo, 2023). However, a research gap remains in understanding how people’s decision-making patterns might influence their use of and reliance on AI.
To address this gap, this paper explores how individuals’ decision-making patterns impact their tendency to rely on seeing knowledge offered by an AI—in our case ChatGPT. As such, we conceptualize individuals’ reliance on AI as their engagement with AI-generated information when AI suggestions are available. When decisions have to be made, who seeks out the suggestions offered by ChatGPT? Who is more susceptible to this information despite the risks? And how are these choices related to individuals’ decision-making patterns?
Our work builds on a well-established decision-making framework proposed by psychologists Irving Janis and Leon Mann (Mann et al., 1997), which has been widely validated with individuals from various countries (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000). Under this framework, decision-making patterns are coping patterns that correspond to situational stress. Adoption of these patterns varies among individuals since people cope with decision-related stress differently, from vigilantly searching for information on their own to deferring decision-making to others. These decision-making patterns can be closely mapped to current interactions with AI systems: the amount of information individuals search during decision-making may impact how much information they need when evaluating AI suggestions; whether individuals defer decision-making to others in human-human interactions could influence their reliance on AI during human-AI interactions.
To explore whether individual decision-making patterns predict a person’s AI-assisted decision-making tendencies, we preregistered [^1] See our preregistration on the Open Science Framework: https://osf.io/z9527/?view_only=ea63631787a74b3c8edd7a8ddc2dbd6b and conducted an online experiment (n=810) in which we asked participants to decide whether statements about food are facts or myths, letting them choose whether they wanted to seek ChatGPT’s decisions and explanations. To increase the stakes, participants were told to put together a pamphlet that should only include factual nutrition statements, and to imagine that the pamphlet will be shared with an organization that focuses on improving children’s health.
Our results show that people differ in their interaction with ChatGPT depending on their individual decision-making pattern. Concretely, the higher a participant’s tendency to be a vigilant decision-maker, the more time they spent looking through ChatGPT’s explanations, underlining their tendency to scrutinize information before deciding. Conversely, the higher a participant’s tendency for buckpassing, the more likely they were to choose to see ChatGPT’s decisions, the less time they spent looking at its explanations for the decision, and the higher their self-reported reliance on ChatGPT. These findings indicate that, in general, buckpassers are less likely to question information provided by an AI, making them more susceptible to misinformation in AI responses than decision-makers who score low on buckpassing or high on vigilance.
Overall, our work contributes (1) new empirical insights into how people differ in their interactions with AI, implying that parts of the population may be more vulnerable to the risks of AI; (2) design implications for AI tools that go beyond one-size-fits-all, including suggestions for leveraging the benefits of AI technologies while mitigating their risks; (3) a new, publicly available dataset (available at the GitHub repository [^2] https://github.com/Mooniem/passing_the_buck_to_AI) with the demographics, decision-making patterns, and AI interactions of 810 participants, which can be used by researchers for replicating and extending our analyses.
2. Related Work
In this section, we first describe prior studies on AI-assisted decision-making to provide some background on common terminology. This section also sets the stage for our subsequent exploration of factors that have been found to influence human-AI interaction—in particular trust and reliance—and how decision-making patterns may be a missing, but important piece in the puzzle.
2.1. Studies of AI-assisted Decision-Making and Terminology
AI-assisted decision-making, also known as human-AI decision-making, broadly refers to scenarios where AI provides suggestions to help humans make decisions (Chen et al., 2023; Lai et al., 2023). To investigate the effectiveness of providing AI suggestions, recent research has examined users’ trust and reliance behaviors such as overall AI reliance, appropriate reliance, over-reliance, and under-reliance. Overall AI reliance measures users’ reliance behaviors in AI-assisted decision-making regardless of the desirability of such reliance. Users’ trust and reliance on systems are often captured via objective measures such as the agreement of an individual’s final decision and an AI’s suggestion in decision-making tasks (Papenmeier et al., 2022; Lu et al., 2024; Cao and Huang, 2022; Cao et al., 2023; Yin et al., 2019; Morrison et al., 2023) or switching from one’s decision to the suggestion of an AI (Yin et al., 2019; Morrison et al., 2023; Cao et al., 2023). While behavioral metrics were also used to quantify individuals’ trust in AI systems (Yin et al., 2019), later researchers pointed out that trust and reliance are distinct concepts with different properties (Lee and See, 2004). Trust is considered to guide yet not necessarily determine reliance behaviors. Thus these behavioral metrics are often used to measure reliance, and separate trust-related questionnaires are used to measure trust in AI systems. Subjective measures such as perceived reliance are also often used in prior studies (Cao and Huang, 2022; Buçinca et al., 2021). Generally, appropriate reliance is desired because it means that users only adopt the AI’s suggestions if they are correct. In contrast, over-reliance (when users adopt the AI suggestions despite it being inaccurate (Buçinca et al., 2021)) and under-reliance (where users fail to adopt the correct AI suggestions (Vasconcelos et al., 2023)) are undesirable. Notably, studies examining these user behaviors often employ a two-stage decision-making workflow in which AI suggestions are integrated into the collaboration process, by automatically presenting the AI suggestions before or after the users make their decisions (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024). As such, the measurement of reliance is often limited to the experimental design that focuses on agreement with AI’s decision or switching. However, to examine whether AI suggestions are effective, it is also critical to understand whether and how people consider them, which might not be universal among individuals. Thus, our study extends the prior behavioral metrics for reliance and takes into account whether and how people choose to see AI suggestions.
2.2. Factors that Influence Trust and Reliance on AI
Recent studies find variance in individuals’ trust in, and reliance on AI suggestions and several factors related to system design and individual characteristics. In terms of system-related factors, evidence from empirical studies identified the effects of design and performance of AI systems such as the accuracy of the AI and whether the AI provides a reason for its decision (Lai and Tan, 2019; Wang and Yin, 2023; Buçinca et al., 2021; Schemmer et al., 2023; Schoeffer et al., 2024). Regarding user characteristics, evidence from survey studies suggests that users’ attitudes towards AI systems vary across demographics, including age, education levels, and knowledge about AI (Gillespie et al., 2023; Ma and Huo, 2023; Araujo et al., 2020; Chong et al., 2022). For example, in a study on individuals’ acceptance of ChatGPT, older individuals are found to be less willing to accept ChatGPT compared to younger individuals (Ma and Huo, 2023). Araujo et al. (2020) find that education levels and knowledge about AI and algorithms are positively associated with users’ perceived usefulness and fairness of AI. In addition, personal traits have also been found to play a role in users’ adoption of suggestions from AI systems: Chong et al. (2022), for instance, find that self-confidence directs individuals’ decisions to accept or reject suggestions from the AI. Focusing on individuals’ Big-Five personality traits (including openness, conscientiousness, agreeableness, neuroticism, and extraversion), Cai et al. (2022) identify that individuals who score high in conscientiousness—a tendency related to self-control and responsibility—have higher trust in AI systems that offer both human-requested and system-initiated suggestions. Faruk et al. (2023) show that students from India and Thailand who score high in openness (being open to new experiences) are more likely to use ChatGPT. Meanwhile, they find negative effects of agreeableness (being cooperative, friendly) and neuroticism (high tendency toward negative feelings) on students’ usage of ChatGPT. Our study extends this line of work by focusing on the effects of individual decision-making patterns in AI-assisted decision-making. While decision-making patterns are found to relate to decision self-esteem and have points of contact with individual differences on personality dimensions such as self-efficacy (Mann et al., 1997), they hold distinct value in human decision-making theory. Unlike personality dimensions that are stable traits, these patterns are individuals’ coping strategies with situational stress, which can be influenced by context, offering value for design interventions that help adapt to individuals’ decision states.
2.3. Decision-Making Theories and Individual Patterns
Scholars from various social science disciplines—such as economics and psychology—have developed different theories explaining the mechanisms behind individuals’ decision-making. The rational choice theory, rooted in economics, posits that individuals calculate the utilities of all options to make decisions. Challenging this proposition, Simon (1955) argues that people often seek satisfactory rather than optimal choices due to their limited capacity to evaluate numerous options. Schwartz et al. (2002) propose that people tend to either have a tendency of maximizing or satisficing: maximizers tend to assess options carefully and choose one that provides maximum benefit later on, while satisficers tend to settle for an acceptable decision. Different from the economists’ perspectives that focus on the utility of decisions, psychologists develop constructs and models of decision-making that relate to motivational processes, cognitive styles, and personality dimensions and traits. The motivational processes refer to decision-making processes in response to certain motivations such as preserving self-image or dealing with negative thoughts and feelings. Example constructs for cognitive styles include need for cognition (the tendency to exert greater cognitive effort to evaluate information) (Cacioppo and Petty, 1982). Mann et al. (1997)’s theory takes into account individuals’ emotions and stress-coping patterns (motivational processes) in their decision-making process. Their four decision-making patterns are derived from a survey study and are described as (i) vigilance (making decisions only after a comprehensive search of information), (ii) hypervigilance (approaching decisions in a hurried and anxious way), (iii) buckpassing (avoiding decisions or deferring to others), and (iv) procrastination (putting off the decision). See Table 1 for an overview and example scale items.
According to Janis and Mann (1977), all four decision-making patterns are in the repertoire of individuals, but people tend to rely on one of these patterns more than on others. Importantly, Mann et al. (1997)’s decision-making patterns have been widely validated with participants from various countries and demographics (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000) and have been shown to be relevant in real-world decision-making scenarios (Brown et al., 2016; Alexander et al., 2017; Kim et al., 2022). Moreover, Mann et al. (1998) reveal that individuals’ reliance on decision-making patterns (especially buckpassing) varies between East Asian and Western cultures. Together these studies suggest that people vary in their approach to decision-making. They also found that individuals’ decision self-esteem is positively correlated with vigilance and negatively with hypervigilance and buckpassing.
Decision-making patterns have been found to affect individuals’ decision outcomes. In the context of economic choices, research shows that when given a time constraint, maximizers tend to browse more options and change their decisions before making the final purchase than satisficers (Chowdhury et al., 2009). Adopting Mann et al. (1997)’s decision-making categories, Kim et al. (2022) found that individuals’ decision-making patterns are linked to their quality of life, including physical and psychological. These findings shed light on the impact of decision-making patterns on individuals’ lives, which could also be manifested in the context of AI-assisted decision-making. Jugovac et al. (2018) has examined how maximizers and satisficers differ in their interactions while presenting AI recommendations; however, they found no effect. Notably, their study design focuses on participants’ interactions with two provided recommendations instead of a design setting where relying on an AI is optional. Our study further explores this line of work in a design setting where AI support is optional, examining how users’ behaviors differ when they are faced with uncertain information and have the option to defer to the AI.
3. Methodology
3.1. Hypotheses
We developed our hypotheses based on the decision-making patterns by Mann et al. (1998) (summarized in Table 1). Importantly, these decision-making patterns emphasize whether people delegate decision-making to others and how they search for information under decision stress. As such, we believe these decision-making patterns are likely predictive of behaviors that we commonly see when humans interact with an AI, such as relying on an AI to seek further information or overrelying on an AI to avoid making one’s own decision.
Table 1. Decision-Making Patterns, Descriptions, and Example Statements from the Melbourne Decision-Making Questionnaire
Decision-Making Pattern Description Example Item
Vigilance This pattern refers to a rigorous information search, such as when a person tries to evaluate different alternatives before making a decision. “I consider how best to carry out the decision.”
Hypervigilance This pattern refers to a frantic way of making decisions to relieve emotional stress. “After a decision is made I spend a lot of time convincing myself it was correct.”
Buckpassing This pattern refers to a tendency to leave decisions to someone else. “I prefer that people who are better informed decide for me.”
Procrastination This pattern refers to a tendency to put off making decisions. “I waste a lot of time on trivial matters before getting to the final decision.”
[…]
Overall, we hypothesize that decision-making patterns could affect individuals’ interactions with AI suggestions. Specifically, by AI suggestions, we refer to the recommended decisions from the AI (AI decisions) and its text-based justifications for its recommendation (AI explanations). Our study imagines scenarios where people encounter information online that may be right or wrong and use an AI tool (e.g., ChatGPT) to assess its credibility. When using an AI, people can seek the AI’s decision and, if needed, the AI’s explanations for the decision. We speculate these AI suggestions are used differently across individuals with varying decision-making patterns, as outlined in Table 2. For vigilant decision-makers who tend to conduct a comprehensive search of information and evaluate it rigorously, we hypothesize that they are more likely to seek the AI’s suggestions for additional information even if they already have an answer. However, as they also tend to be more confident in their own decisions (Mann et al., 1998), they might not necessarily prioritize the AI’s decision over their own.
People who exhibit hypervigilant decision-making behaviors tend to search for information frantically, but not rigorously, before arriving at a decision. Thus, we hypothesize they will quickly seek information from the AI (i.e., see AI decision and explanations) without spending much time evaluating the information. As they experience high emotional stress and are less confident in their decisions (Mann et al., 1998), they would be more likely to rely on the AI’s suggestions to reach their decisions.
We further hypothesize that people who exhibit a buckpassing tendency will seek out the AI’s suggestions to avoid making any decisions themselves, leading to a high reliance on AI. As a result, they would not consider AI’s suggestions as much, leading to less time spent on AI explanations.
Table 2. We develop our hypotheses based on individuals’ decision-making patterns, focusing on their behavioral characteristics related to information search.
Decision Making Pattern Hypotheses Notation
The higher people score on vigilance, the more likely they are to choose to see AI decision. H1a
the more time they spend seeing AI explanations. H1b
the lower they report their reliance on AI to be. H1c
The higher people score on hypervigilance, the more likely they are to choose to see AI decision. H2a
the less time they spend seeing AI explanations. H2b
the higher they report their reliance on AI to be. H2c
The higher people score on buckpassing, the more likely they are to choose to see AI decision. H3a
the less time they spend seeing AI explanations. H3b
the higher they report their reliance on AI to be. H3c
We did not include procrastination in our hypotheses because this would require leaving participants with potentially infinite time, which we did not deem feasible. However, procrastination has been shown to be correlated with behaviors or buckpassing and hypervigilance (Bouckenooghe et al., 2007), and all three decision-making patterns are considered forms of defensive avoidance (Janis and Mann, 1977).
3.2. Experiment Design
To examine our hypotheses, we designed an online study (preregistered at link) that includes (i) a questionnaire and (ii) a task requiring participants to decide on the factuality of a series of nutrition statements while having the option to seek out AI suggestions (AI decisions and AI explanations). We generated AI decisions and explanations using ChatGPT before the study (rather than having participants interact with ChatGPT live) to ensure consistency of response quality. The study was approved by our Institutional Review Board and launched in English on the volunteer-based online study platform LabintheWild [^3] https://labinthewild.org/. We encouraged honest participation by informing participants in advance that they would receive personalized result—their own decision-making patterns and performance on evaluation food facts and myths—at the end of the study.
Nutrition Statements for the Main Task: We used 30 nutrition statements and their accuracy (i.e., fact or myth) from Florença et al. (2021) who curated a list of popular food (mis)-conceptions from online sources. In their study with 503 participants, several of these statements were inaccurately classified as facts or myths by participants, and even by participants working in areas related to nutrition, suggesting that the statements have a range of difficulty. Among the 30 statements, 9 nutrition statements are facts and 21 are myths. One example false statement (a myth) is “Pregnant women should be eating for two”, while an example true statement (a fact) is “Dairy products should be consumed in between two and three portions per day”. The full list of statements with fact/myth labels is in the Supplementary Materials.
AI Suggestions (AI Decisions and Explanations): We generated AI decisions and explanations for each nutrition statement using ChatGPT [^4] https://chat.openai.com/ in July 2023, prompting ChatGPT in the format of “[Nutrition Statement]. Is this fact or myth?” After the model generated an answer to the prompt, we asked it to shorten its answer to less than four sentences by prompting “Rewrite it in less than 4 sentences.” Based on nutrition statements from Florença et al. (2021), ChatGPT produced incorrect answers for three out of the 30 statements (two false facts and one false myths). We used ChatGPT’s decision whether the statement is a fact or a myth as the AI Decision and the rest of the information it provided as the AI Explanation.
We then chose 21 accurate AI suggestions (each comprising a decision and explanation), the 3 inaccurate AI suggestions that ChatGPT provided, and further inverted 6 suggestions to become false, resulting in 21 accurate and 9 inaccurate AI suggestions. From these two pools of statements, we randomly sampled 6 statements with accurate AI suggestions and 4 with inaccurate AI suggestions. We included more accurate than inaccurate AI decisions in our experiment to be faithful to the real-world performance of AI models (i.e., ChatGPT provided accurate responses to most of the 30 nutrition statements). To generate inaccurate AI explanations for these 6 statements, we prompted ChatGPT with ‘‘What are some arguments people would use to support this statement as a fact/myth: [Nutrition Statement].” and also asked it to shorten its answer to have fewer than four sentences.
3.3. Study Procedure
Our online study, directed at intrinsically motivated volunteer participants, was advertised with the question “What is your decision-making style?” Participants were able to receive personalized results at the end of the study. As they entered the study page, we explained the goal of the experiment and that they would be asked to judge a set of nutrition statements. Participants could proceed to the experiment only if they consented. Participants were then asked to fill out a demographic questionnaire including questions about their age, gender, education levels, and country. We also inquired about their knowledge of nutrition (“How would you describe your knowledge of nutrition compared to the general public?” with the answer options being: “Very limited”, “Limited”, “Average”, “Above average”, and “Expert”) and assessed their perception of the accuracy of AI-generated information (“To what extent do you believe that conversational AI (e.g. ChatGPT) provides correct information?” with answer options “I don’t know what ChatGPT is”, “It never provides anything correct”, “It provides correct information sometimes”, “It provides correct information usually”, and “It always provides correct information.”).
Next, participants were asked to complete the Melbourne Decision-Making Questionnaire (MDMQ) (Mann et al., 1997) (see Supplementary Materials), with 10 questions being shown before the main tasks and the remaining 7 questions being shown after the main task (all in random order) to reduce the perceived questionnaire length.
Figure 1. The study interface and workflow for each statement. Participants first read one statement at a time. They could then decide whether the statement is a fact or myth ( a). If participants decided at this point, they could move on to the next statement. Alternatively, participants could choose to reveal the AI’s decision ( 1) showing ChatGPT’s decision ( b). Participants could then decide to choose fact or myth or use the AI decision. They could also choose to see AI’s explanation to reveal the AI explanation before making their decision ( 2).
This figure illustrates the interaction flow of the main task for participants. The figure indicates two possible interaction flows for participants when they are evaluating food facts or myths. One of the interaction flows is that participant can choose to provide their answer based on their own knowlege and select the correponding answer. The other workflow is particpants can choose to select See AI decision button to see AI’s decision for the evaluation of statement. After the AI’s decision is revealed, participants would be shown a button that says ”Show me the AI’s explanation”. At this point, they can either skip the explanation and just make their final evaluation, or click the button to reveal AI’s text-based explanations.
Main task: The main task showed participants 10 nutritional statements (6 accurate and 4 inaccurate decisions, drawn from the larger set of 30 statements using stratified random sampling). For each statement, participants were asked to assess whether it was a fact or myth while being provided options to use AI assistance (see Figure 1). To encourage participants to make cautious decisions about these statements, we introduced the following scenario:
Imagine your manager asks you to put together a nutrition pamphlet to share with an organization that focuses on improving children’s health. To do so, you will need to evaluate 10 nutrition-related statements and determine if they are facts or myths. You will have access to an AI that can assist you with these assessments. However, please keep in mind that the AI’s suggestions may not always be entirely accurate or definitive. Whether to use the AI’s decision is completely up to you.
We mentioned the limited performance of AI assistance to ensure participants knew that not all information presented to them was factually correct, in line with current state-of-the-art conversational AI systems which do not always output accurate information (Deiana et al., 2023). Informing participants about the imperfection of decision aid is often used in prior work assessing users’ trust and reliance on automated support systems (Dzindolet et al., 2003). While it might prime participants to refrain from seeking AI suggestions, it could also encourage participants to critically evaluate AI suggestions rather than accept them blindly.
If participants chose “I’d like to see the AI decision,” they were provided two additional options: to see an AI explanation or to immediately accept the AI decision (without seeing an explanation). This design was motivated by our hypotheses: see AI decision is designed to observe whether users prefer to refer to the AI’s decision before making their own. Subsequently, we offered the option to see AI explanations to investigate whether some users might want to further investigate and seek additional details.
After an AI decision was revealed, participants could select Use the AI decision to indicate they accept the AI decision; however, this design revealed its limitations as participants barely selected this option based on our analysis of results which we discuss later. For each statement, we recorded the time they took to evaluate the AI’s decision and explanations before final decision-making.
At the end of the study, participants were asked to assess their reliance on AI during the study via the question “How much did you rely on the AI in your decisions for food facts and myths?”. Participants could select from a five-point scale with 1 being ‘Didn’t rely on it at all’ and 5 being ‘I relied on it all the time’.
The experiment ended with a page showing participants their decision-making patterns, how many nutrition statements they judged correctly versus incorrectly, and which statements they got wrong. We further debriefed participants by stating that we used ChatGPT for the AI-generated content during the experiment and that only some of the nutrition statements and explanations were correct.
3.4. Metrics
We operationalize individuals’ interactions with AI suggestions via a combination of clicking behaviors, reading time, and self-reported reliance:
  • See AI Decision is a binary variable that indicates whether participants select the option “I’d like to see the AI’s decision” ( 1 in Figure 1) for each statement. This variable measures whether participants seek more information from the AI during the task. For each participant, we also calculated the percentage of statements they see AI decisions (Frequency of See AI Decision).
  • See AI Explanations is a binary variable indicating whether participants selected the option “Show me the AI explanation” ( 2 in Figure 1) for each statement. This measures whether people seek more information from the AI in the task after seeing the AI decision. For each participant, we also calculated the percentage of statements they see explanations out of the 10 statements they see (Frequency of See AI Explanations).
  • Time Spent Seeing AI Explanations is a continuous variable that measures the time it takes participants to select a final evaluation of a statement after they choose to see the AI explanation. We only examined this variable for participants who select a response after they click “See AI Explanation”.
  • Perceived AI Reliance measures individuals’ self-reported reliance on AI assistance in the study based on their responses to the question “How much did you rely on the AI in your decisions for food facts and myths?” We coded levels 1 to 5 (1 for “Didn’t rely on it at all”, 5 for “I relied on it all the time”). We adopted a one-item measurement approach to examine individuals’ perception-based reliance on AI based on prior study design (Cao and Huang, 2022). The item was adapted to focus on our nutrition task, assessing participants’ reliance explicitly during the study.
We examine the effects of decision-making patterns while considering the demographic covariates (age, gender, education levels) as well as the following variables:
  • Vigilance, Hypervigilance, Buckpassing: Based on participants’ answers to the Melbourne Decision Making Questionnaire (MDMQ) (Mann et al., 1997) on a 3-point scale (“True for me” (score 2), “Sometimes true for me” (score 1), “Not true for me” (score 0)), we obtained the total score for each of their decision-making dimensions. Each decision-making pattern has its own subscale, consisting of a unique set of statements distinct from those used in the other subscales. Therefore, while we excluded procrastination in our hypotheses and did not administer that portion of the MDMQ, this should not impact the validity of measuring buckpassing, vigilance, and hypervigilance. The maximum score is 10 for hypervigilance, and 12 for both vigilance and buckpassing. The vigilance and buckpassing subscales consist of 6 items each, while the hypervigilance subscale consists of 5 items.
  • Domain Knowledge is based on participants’ self-reported knowledge level of nutrition compared to the average public. We categorized participants into three groups (“Below Average”, “Average”, “Above Average”) due to the imbalance of individuals among the five subgroups (“Very Limited” (41), “Limited” (92), “Average”(351), “Above Average”(300), “Expert” (27)). When participants reported their knowledge levels as “Above Average” or “Expert”, their domain knowledge was coded as “Above Average”. We categorized “Very Limited” and “Limited” to be “Below Average”. Note that we treated this variable as nominal since we were uncertain about whether the increase of likelihood to see AI suggestions and perceived AI reliance would be the same from “Below Average” to “Average” compared to from “Average” to “Above Average”.
  • Perception of AI of participants were coded as “High” and “Not high” based on their perception of the accuracy of conversational AI before participating in the experiment. If they selected “It provides correct information usually” or “It always provides correct information”, it was coded as “High”, otherwise it was coded as ”Not high”.
3.5. Participants
Participants in our study were 917 volunteers recruited via the online study platform LabintheWild between August 2023 and April 2024. The platform was chosen to allow us to recruit diverse participants of various ages and education levels to ensure a broad spread of decision-making styles. Participation on LabintheWild is open to anyone without signing up; we therefore obtained a waiver of parental consent from our IRB so that minors were able to take our study. To reduce the risk of participants mistaking factually incorrect nutrition information as correct, the study instructions clearly stated that not all nutrition statements are factual and that the goal is to test how much participants know about nutrition. As part of this debrief, participants were also told which of the nutrition statements they wrongfully assumed to be true at the end of the experiment.
We excluded any participants who did not complete the study or affirmed they had taken the study before. To reduce the risk of including participants who were not truthfully responding to the task (i.e., those exhibiting satisficing behavior (Kapelner and Chandler, 2010)), we also excluded participants who consistently selected either the option “True for me” or “Not true for me” for all questions in the decision-making questionnaire. After examining the distribution of participants’ total time spent completing the study (included in the Appendix), we decided to remove participants whose study completion times were outliers—either extremely short or extremely long—and unlikely to represent a serious attempt at completing the tasks. On average, participants took 9 minutes to finish the study (Median = 468 seconds, SD= 297 seconds). Thus, we eliminated participants who completed the study in less than 240 seconds minutes (N=18) or more than 1200 seconds (N=54), accounting for participants who did not seek AI suggestions and those who reviewed all suggestions. In total, we excluded 107 participants, resulting in a final sample of size 810.
The final set of participants reported being from 69 countries, with the majority from the United States (44%), followed by the United Kingdom (9%), Canada (8%), India (5%), Germany (3%), and Australia (3%, all others $\leq 3\%$). Half of our participants reported pursuing or having obtained a college education, approximately 29% of participants reported pursuing or obtaining a high school education; with the rest reporting to pursue or having obtained a graduate education. Our study participants were between 12 to 90 years of age ($M=31$, $SD=14$). A detailed breakdown of demographics for participants is shown in Table 3.
Table 3. Sample sizes of study participants across different demographics in terms of participants’ education levels, gender, age group, self-reported knowledge of nutrition, and their perception of the accuracy of AI-generated information.
N
Overall
   Overall 810
Education
   ¡=High School Education 231
   Pursuing or Have obtained college education 403
   Pursuing or Have obtained graduate education 176
Gender
   Female 409
   Male 363
   Non-binary/No-disclosure 38
Age Group
   Less than 18 106
   18-24 264
   25-34 189
   35-44 107
   45-55 82
   Above 55 62
Nutrition Knowledge
   Very limited 41
   Limited 92
   Average 351
   Above Average 300
   Expert 26
Perception of AI
   I don’t know what conversational AI (ChatGPT) is. 64
   It never provides anything correct. 8
   It provides correct information sometimes. 346
   It provides correct information usually. 378
   It always provides correct information. 14
This table shows the sample sizes of study participants across different demographics, including their education levels, gender, age group, self-reported knowledge of nutrition, and their perception of the accuracy of AI-generated information. The table contains columns for the different categories and the corresponding sample sizes for each.
4. Analyses and Results
4.1. Descriptive Statistics
This section presents our sample characteristics and key variables, as a form of robustness check in line with prior work Mann et al. (1998) and Florença et al. (2021). Then we present detailed analysis results on (1) Effect of Decision-Making Patterns on Seeing AI Decisions (H1a, H2a, H3a), Effect of Decision-Making Patterns on Seeing AI Explanations (H1b, H2b, H3b), and Effects of Decision-making Patterns on Perceived AI Reliance (H1c, H2c, H3c).
Variance in Decision-Making Patterns and Measurement Reliability
The full summary statistics of decision-making patterns for the different demographics are shown in the Appendix (Table 9). On average, our study participants scored 9.41 ($SD=2.31$) out of 12 on vigilance, 4.25 ($SD=2.61$) out of 10 on hypervigilance, and 4.56 ($SD=3.03$) out of 12 on buckpassing. Each subscale reaches a satisfactory level of internal reliability. Cronbach’s alphas for each subscale is 0.74 (vigilance), 0.81 (buckpassing), and 0.74 (hypervigilance).  [^5] Cronbach’s alpha measures internal consistency and reliability of a scale, ranging from 0 to 1. Values from 0.70-0.90 are recommended to be acceptable values (Tavakol and Dennick, 2011). Cronbach’s alphas were 0.80,0.87,0.74 from Mann et al. (1998). The mean and standard deviation of our measures fell within a similar range compared to those observed in Mann et al. (1998) across different countries. Participants’ hypervigilance had significantly positive correlation with their measure of buckpassing ($r=0.59,p<0.001$), which aligns with prior studies (Filipe et al., 2020; Mann et al., 1997). Additional factor analysis confirmed that these three different decision-making patterns were identified with the same corresponding question items from Mann et al. (1998). Details regarding factor loadings and a corresponding diagram are included in the Appendix. These similarities suggest the reliability of our measure of decision-making patterns and the robustness of our samples.
Variance in Participants’ Frequency of Seeing AI Decisions and Explanations
We first verified if there was variance in participants’ interactions with AI suggestions before we identified the respective underlying factors. We found that the frequency of seeing AI suggestions varied across participants, with the average participant choosing to see the AI’s decision $30\%$ of the time ($M=30\%,SD=26\%,Min=0\%,Max=100\%$) and the AI’s explanations $23\%$ of the time ($M=23\%,SD=23\%,Min=0\%,Max=100\%$).
Participants’ Overall Performance
On average, participants accurately evaluated 70% of the statements ($M=70\%,SD=14\%$). Participants who scored high in buckpassing (M = $68\%$) had a slightly lower accuracy than those who scored average (M = $70\%$) and low in buckpassing (M = $71\%$). When AI decisions were incorrect, participants scored much lower in accuracy ($M=23\%$) when they chose to see AI decisions compared to times when they did not ($M=57\%$). When AI decisions were correct, participants scored slightly higher in accuracy when they chose to see AI decisions ($M=90\%$) compared to times when they didn’t ($M=84\%$). A detailed breakdown of accuracy across different conditions and buckpassing level is included in the appendix (Figure 10). Performance did not significantly differ across the different decision-making styles. The performance of our participants on each statement was significantly correlated with participants in Florença et al. (2021) ($r=0.58,p<0.001$), indicating that most of our participants exhibit a similar knowledge pattern regarding common food myths and facts.
Agreement with AI suggestions
To examine participants’ decision behaviors after they see AI decisions, we measured their agreement rate. Proportion of agreement with AI suggestions is often used as a behavioral metric of reliance in prior studies where AI suggestions are automatically shown (Lu et al., 2024; Cao and Huang, 2022; Cao et al., 2023). However, since our study design provided AI suggestions on demand, we measured participants’ conditional agreement rate conditioned on whether they chose to see AI decisions. Participants agreed with AI decisions on about 83% of statements when they chose to see them, compared to 67% when they did not. Among all participants, around 49% (395 out of 810) of participants agreed with AI every time they chose to see AI decisions. We also visualized the distribution of participants’ individual conditional agreement rates in the Appendix (Figure 9).
Figure 2. Overview of Frequency of See AI Decisions with Different Levels of Buckpassing: Participants with varying levels of buckpassing show differences in how often they view AI decisions. “Low” and “High” buckpassing refer to scores that are more than one standard deviation below and above the average, respectively. “Average” buckpassing refers to scores that are within one standard deviation of the average. Each bar represents the average frequency, with error bars indicating the confidence intervals. On average, those who score low in buckpassing view AI decisions 25% of the time, while those who score high view them 33% of the time.
This figure is a bar chart displaying the relationship between buckpassing (a decision-making style) and the frequency of users choosing to ”See AI Decisions” across three buckpassing groups: Low, Average, and High. On the y-axis is the average frequency of selecting ”See AI Decisions” and the error bar. The scale ranges from 0.0 to 0.4. On the x-axis is the buckpassing levels divided into three categories: low, average, and high. Low level includes participants whose buckpassing score is 1 standard deviation below the mean (marked in red). Average level includes participants with an average buckpassing score. High level includes participants whose buckpassing score is 1 standard deviation above the mean. Sample size for each group are displayed above the bars: Low buckpassing group has n = 139. Average buckpassing group has n = 540. High buckpassing group has n = 132. The heights of the bars show that participants in the Low and High buckpassing groups selected ”See AI Decisions” less frequently than those in the Average group. The error bars are 95% confidence intervals for the mean frequencies. The Low buckpassing group has a slightly lower frequency with a narrow confidence interval, while the Average and High groups have higher frequencies with overlapping confidence intervals.
4.2. Effect of Decision-Making Patterns on Seeing AI Decision
To examine the effect of decision-making patterns on individuals’ likelihood to see AI decisions (H1a, H2a, H3a), we fit a series of mixed-effects logistic regression models for our binary dependent variables (i.e., whether a participant decided to see the AI decision and the AI explanation for each statement). We included participants’ individual decision-making scores (for vigilance, hypervigilance, buckpassing) as fixed effects, with participant and nutrition statement as random effects (Model 1 in Table 4 ) to account for variation across nutrition statements and individual differences. Neither hypervigilance nor vigilance showed a significant effect, suggesting that H1a and H2a are not supported. Buckpassing had a significant positive effect on the likelihood of people seeing AI decisions ($\beta=0.062,p<0.05$), supporting H3a. We also show an overall frequency of seeing AI decisions across different levels of buckpassing among our participants in Figure 2.
Table 4. Mixed-effects logistic regression results predicting participants’ tendency to seek AI suggestions (See AI Decision and See AI Explanations). Regression analysis with only variables related to decision-making patterns indicates a statistically significant effect of buckpassing on one’s likelihood to see AI’s decision. When including demographic variables, the regression analysis shows that participants’ education levels, self-reported domain knowledge, and perception of AI-generated information have significant effects on participants’ tendency to seek AI suggestions. Reference groups are less than college education, domain knowledge (Above average), and perception of AI (Not high).
Dependent variable:
See AI Decision See AI Explanations
Model 1 Model 2 Model 3
Coef. (S.E.) Coef. (S.E.) Coef. (S.E.)
Buckpassing 0.062∗∗ (0.025) 0.048 (0.025) 0.032 (0.026)
Hypervigilance -0.001 (0.029) -0.021 (0.029) -0.029 (0.030)
Vigilance 0.015 (0.027) 0.014 (0.026) -0.007 (0.027)
Age -0.015∗∗∗ (0.005) -0.016∗∗∗ (0.005)
Pursuing or Have obtained college education 0.071 (0.149) 0.104 (0.153)
Pursuing or Have obtained graduate education 0.624∗∗∗ (0.185) 0.631∗∗∗ (0.190)
Domain_Knowledge (Average) 0.594∗∗∗ (0.136) 0.406∗∗∗ (0.139)
Domain_Knowledge (Below Average) 0.500∗∗∗ (0.187) 0.108 (0.193)
Perception of AI (High) 0.256∗∗ (0.121) 0.080 (0.124)
Constant -1.697∗∗∗ (0.315) -1.695∗∗∗ (0.387) -1.704∗∗∗ (0.399)
Observations 8,100 8,100 8,100
Akaike Inf. Crit. 8,411.871 8,379.897 7,455.465
Marginal $R^{2}$ / Conditional $R^{2}$ 0.006 / 0.470 0.033 / 0.469 0.020 / 0.469
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01
[…]
Additionally, we built a model (Model 2) that controls for demographic covariates since prior research suggests that individuals’ demographics factors such as age, education levels, domain knowedge (Kim et al., 2023; Ma and Huo, 2023; Araujo et al., 2020) influence their trust and reliance on AI. The details of this model are summarized in Table 4. The positive effect of buckpassing remains marginally significant after controlling for participants’ age, education, domain knowledge, and perception of an AI ($\beta=0.048,p=0.06$).
4.3. Effect of Decision-Making Patterns on Seeing AI Explanations
To test H1b, H2b, and H3b, we first examined whether participants chose to see AI explanations, which was true for 548 out of 811 participants. Per Model 3 in Table 4, there is no significant effect of decision-making patterns on individuals’ tendency to see AI explanations. We included the results of the model with only decision-making measures as fixed effects in the Supplemental Materials.
Next, we examined the effect of decision-making patterns on how much time people spend reviewing AI explanations, if they choose to see them. On average, participants spent 14 seconds seeing AI explanations ($M=14.34,SD=10.73$). We used a linear mixed-effects regression model, considering that each participant could have their own reading speed and each statement had different lengths of AI explanations. Per Table 5, buckpassing had a statistically significant negative effect on individuals’ time spent evaluating AI explanations ($\beta=-0.40,p<0.05$). For one additional score of increase in buckpassing dimension, the time users spend on seeing AI explanations decreased by 0.4 seconds. This suggests individuals who score higher in buckpassing tend to spend less time evaluating AI explanations before they make their final decisions. On the contrary, vigilance had a statistically significant positive effect on participants’ time spent evaluating AI explanations ($\beta=0.53,p<0.05$). For one score of increase in vigilance dimension, the time users spent on reading explanations increased by 0.53 seconds. These results support our hypotheses H1b and H3b, but not H2b.
Table 5. Linear mixed model predicting individuals’ time spent evaluating AI explanations: participants’ buckpassing and vigilance tendency significantly affect the time they spent reading AI’s explanations if they chose to see them.
Dependent variable:
Time Spent Seeing AI Explanation (Seconds)
Buckpassing -0.401∗∗ (0.197)
Vigilance 0.528∗∗ (0.206)
Hypervigilance 0.291 (0.223)
Age 0.054 (0.038)
Domain_Knowledge (Average) -0.162 (1.035)
Domain_Knowledge (Below Average) 4.173∗∗∗ (1.454)
Pursuing or Have obtained college education $0.026$ (1.160)
Pursuing or Have obtained graduate education -0.066 (1.431)
Perception of AI (High) 0.885 (0.931)
Constant 6.950∗∗ (2.864)
Observations 1,542
Marginal R2 / Conditional R2 0.023 / 0.232
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01
This image shows the results of a linear mixed model predicting individuals’ time spent evaluating AI explanations. The table displays the coefficient (standard error) for each independent variable.The independent variables include Buckpassing, Vigilance, Hypervigilance, Age, Domain_Knowledge (Average), Domain_Knowledge (Below Average), Pursuing or Have obtained college education, Pursuing or Have obtained graduate education, Perception of AI (High), and Constant. The table also provides the number of observations and the Marginal R2 / Conditional R2 values. The results indicate that participants’ buckpassing and vigilance tendency significantly affect the time they spent reading AI’s explanations if they chose to see them. Specifically, higher buckpassing is associated with less time spent evaluating AI explanations, while higher vigilance is associated with more time spent evaluating AI explanations
Table 6. Ordinal Logistic Regression Model for Perceived Reliance on AI (reference levels are high school education, domain knowledge-above average, perception of AI-low). Significance levels:p¡0.001***, p¡0.05**, p¡0.1*
Predictors Odds Ratios CI
Buckpassing 1.09∗∗ 1.03 – 1.15
Hypervigilance 0.97 0.91 – 1.03
Vigilance 0.98 0.93 – 1.04
Pursuing or Have obtained college education 1.21 0.89 – 1.64
Pursuing or Have obtained graduate education 1.60 1.08 – 2.37
Age 0.99 0.98 – 1.00
Domain Knowledge (Average) 1.86∗∗∗ 1.40 – 2.47
Domain Knowledge (Below Average) 1.46 0.98 – 2.17
Perception of AI (High) 1.41∗∗ 1.09 – 1.82
Observations 810
$R^{2}$ Nagelkerke 0.056
[…]
4.4. Effects of Decision-making Patterns on Perceived AI Reliance
To test H1c, H2c, and H3c, we constructed an ordinal logistic regression model while controlling for participants’ demographic covariates including age, education level, domain knowledge, and perception of AI. Ordinal logistic regression is often used to predict dependent variables that can be ordered in a natural way such as mild, moderate, severe (Harrell, 2015). Per Table 6, our result indicates the effect of buckpassing on perceived AI reliance was statistically significant ($OR=1.09,95\%\text{ CI}[1.03,1.15],p<0.05$). For one score of increase in buckpassing, the odds of having more self-perceived reliance on the AI increases by 9%, suggesting that people who score high in buckpassing are more likely to rely on AI when making their decisions. Figure 3 shows the distribution of response in terms of level of reliance related to individuals’ buckpassing tendency. Noteworthily, 12% of people who received a low buckpassing score reported to rely on the AI “all the time” (rating of 5) or most of the time (rating of 4), while 20% of those who received a high buckpassing score reported the same.
No effect of vigilance or hypervigilance was observed. Hypotheses H1c and H2c are therefore not supported, while H3c is supported.
Figure 3. Perceived AI Reliance Variations Across Different Levels of Buckpassing: individuals who scored high (one standard deviation above average among participants) in buckpassing reported a higher reliance on AI compared to those who scored low (one standard deviation below average) in buckpassing.
[…]
4.5. Who tends to score high in buckpassing?
Figure 4. Examining the demographic distribution of buckpassing tendency: participants who have a high school education or below score higher in buckpassing (23% in high buckpassing level); participants who are aged under 18 and in their early 20s also tend to score higher in buckpassing. The y-axis is the proportion of participants in each buckpassing group.
Our analysis indicates that a buckpassing tendency significantly influences participants’ use of AI suggestions and reliance on AI. Thus we further explored the demographic distribution of buckpassing scores among our participants to gain insights into which demographic groups exhibit a higher buckpassing tendency.
Figure 4 shows the distribution of buckpassing levels among different education levels and age groups. Age is treated as a categorical variable in this descriptive analysis to obtain a more granular understanding of age group differences. Teenagers and young adults in our sample tended to score higher in buckpassing compared to participants who were aged above 35. Participants who obtain a lower level of education (high school education or below) also exhibited a higher buckpassing tendency compared to those with a higher level of education (pursuing or having obtained college or graduate education).
5. Discussion
Individuals’ decision-making patterns have been shown to influence their decision outcomes and quality of life (Kim et al., 2022; Deniz, 2006), underlining the importance of understanding how people vary in making decisions. However, despite the increasing role of AI in decision-making contexts, how individual cognitive decision-making processes influence the use of AI technology remains unknown. Our study offers evidence of this influence; we discuss our findings and consider their implications for designing AI systems and explainable AI research below.
Decision-making Patterns as Important Factors in AI-assisted Decision-making
In this work, we extend the current understanding of individuals’ interactions with AI-generated information and factors behind these interactions in a setting where AI assistance is available but optional. Consistent with our hypotheses, we found that people who tend to defer decision-making patterns to others (score high in buckpassing) are more likely to seek out suggestions from an AI and report relying on it, though they spend less time reading AI’s explanations. While the difference can be described as a small to medium effect—those who score low on buckpassing viewed the AI’s decisions 25% of the time versus 33% for those who scored high on buckpassing—the effect of this difference can be profound when AI decision-making tools are frequently used and the performance of AI tools is not high. As we observed in our data, participants scored lower in accuracy when exposed to inaccurate AI decisions and higher when exposed to correct AI decisions. The difference may also gain importance with the number of decisions to be made as decision fatigue sets in (Chen et al., 2023) or as task difficulty or uncertainty increases (Vasconcelos et al., 2023; Salimzadeh et al., 2024), leading to a higher likelihood of overreliance.
We encourage future work to examine these factors in various kinds of human-AI interactions. In addition, prior HCI research shows that individuals’ trust and reliance vary by their age, education levels, and perception of AI (Ma and Huo, 2023; Araujo et al., 2020). Our work extends this by showing that individuals’ buckpassing tendency has a marginal effect on AI use when accounting for these demographic covariates. While the significance of the effect decreases after controlling these factors, this is likely due to the correlation between these demographic factors and decision-making patterns (see Appendix Table 10 for the correlation matrix).
What might explain these findings is the stress level and confidence people experience while making decisions. Mann et al. (1997)’s work has indicated that these decision-making tendencies can be seen as different kinds of coping mechanisms during decision conflict. Individuals with a strong buckpassing tendency often experience high psychological stress during decision-making tasks, while vigilant decision-makers experience a lower level of stress (Mann et al., 1997). Experiencing high psychological stress while making decisions might propel people to seek out an immediate relief by turning to AI suggestions. Mann et al. (1998) also identified that individuals who score high in buckpassing tend to have lower confidence in their own decision-making. This could potentially propel participants in our study to seek AI’s suggestions when they were not confident about making the decisions themselves.
Interestingly, we did not observe any effect of vigilance or hypervigilance on participants’ tendency to see AI suggestions. This might be due to the low variance of participants on vigilance ($SD=2.31$) and hypervigilance ($SD=2.61$) compared to buckpassing pattern ($3.03$). However, we observed that participants who scored high in vigilance spent significantly more time seeing AI explanations. This suggests that increased vigilance goes hand in hand with careful reading of any explanations people are provided, likely as a tool for evaluating whether the AI’s decision can be reasonably trusted.
Benefits and Risks of AI-assisted Decision-making for People with Different Decision-Making Patterns
AI-assisted decision-making could yield both positive and negative outcomes for people with a tendency for buckpassing. On the one hand, the rise of AI systems could ease the burden of information seeking and processing for individuals to make effective decisions, if the performance of these systems is reliable. As evidenced by prior work, making decisions with the help of AI could lead to more accurate decisions when the AI has high advice accuracy and quality (Vodrahalli et al., 2021; Tschandl et al., 2020) (though there are some exceptions where this was found not to be the case (Vaccaro and Waldo, 2019; Jacobs et al., 2021)). AI assistance can also lower mental demand (Buçinca et al., 2021). With highly reliable AI systems, AI suggestions could help relieve the decision-making stress experienced by people with a buckpassing tendency. On the other hand, given the current risks of AI hallucinations, where AI generates unfaithful or nonfactual content (Bubeck et al., 2023), and AI overconfidence (Zhou et al., 2024), where AI outputs sound confident yet are inaccurate (Zhou et al., 2024), buckpassers are more susceptible to inaccurate information generated by AI, potentially resulting in misleading decisions and negative decision outcomes.
The benefits and risks of AI-assisted decision-making might be exacerbated for some people and groups, as our study and other research show that decision-making patterns vary across demographics. Consistent with prior findings (Filipe et al., 2020), our study suggests that buckpassing tendency is more prevalent among younger people (especially individuals under 18) and people with high school education or below. This warrants greater intervention to guide these populations in their usage of AI systems. Prior work from Mann et al. (1998) indicates that people from East Asian countries (e.g., Japan, Hong Kong) exhibit a higher buckpassing tendency compared to people in Western countries (e.g., USA, Australia, New Zealand). East Asians have also been found to have a lower decision-making self-esteem, which is generally in line with a preference for making decisions in a group (Mann et al., 1998). These demographic groups may seek out information from the AI more often, benefiting more from AI systems that have high accuracy but also facing greater risks of using inaccurate AI-generated information.
5.1. Implications for Designing AI Systems and Explainable AI Research
The variations in people’s decision-making patterns, and in particular in people’s tendency to be vigilant or pass the buck warrant more caution in the design and implementation of AI systems. How do we cultivate appropriate reliance on AI systems given these decision-making patterns? Prior work has investigated several interventions that could help mitigate inappropriate reliance, which we consider together with the implications of our findings below.
Account for the differential effects of cognitive forcing functions
Our findings showed that those prone to deferring decisions to others are also more likely to seek out AI suggestions. The relatively short time they spent looking at the AI explanations, compared to participants who scored low in buckpassing, further suggests that buckpassers may not engage as strongly with this information; the immediate access to the AI suggestions (as is most common in human-AI studies and AI decision-support systems in various contexts (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024)) seems to be an easy way to lower cognitive effort, which may lead to overreliance on the AI. Cognitive forcing functions—like presenting AI suggestions on demand or forcing people to make their own decisions before seeing an AI’s suggestions—are effective strategies researchers have identified in reducing overreliance (Buçinca et al., 2020). Yet it remains unclear how individuals differ in interacting with each kind of cognitive forcing function. Our results suggest that decision-making patterns also influence individuals’ engagement with AI suggestions when they are presented on demand. In addition, since buckpassers often defer decision-making to others due to experienced stress, implementing cognitive forcing functions may lead to more mental demand and stress in decision-making. Thus, designers should prioritize relatively low-effort cognitive forcing functions and couple them with motivational nudges (Caraban et al., 2019; Vasconcelos et al., 2023) (e.g., by reminding the potential consequences of overreliance behaviors (Bo et al., 2025)) to encourage active engagement among this group.
Adopt presentation formats of AI explanations that ease cognitive load and decision stress.
Our study found that buckpassers tend to spend less, and vigilant people more time, when reviewing AI explanations, posing challenges to the design of AI explanations. The format of AI explanations in our study was text-based and lengthy. This format might induce further cognitive load and stress for buckpassers, resulting in their reduced time investment in AI explanations. Prior research found that if decisions are made under high cognitive load, they can lead to higher overreliance on AI (Zhang et al., 2024). Thus, similar to cognitive forcing functions, identifying presentation formats of AI suggestions that reduce cognitive load and decision stress may encourage buckpassers to evaluate AI explanations carefully. Chen et al. (2023) find that for individuals who rely on their intuition in decision-making, providing example-based scenarios as AI explanations can effectively calibrate appropriate reliance on AI. Psychology research on cognitive load theory (CLT) and e-learning has identified various information presentation techniques to reduce individuals’ cognitive load (Kirschner, 2002; Van Merrienboer and Ayres, 2005; Sweller, 2011). Mousavi et al. (1995) found that presenting information in both audio and visual format leads to less cognitive load than in visual format alone. As such, designers could consider showing AI explanations to users in both visual and audio formats to lower their cognitive load. Redundant information is also shown to increase individuals’ cognitive load (Sweller, 2011). However, experts and novices have different perceptions of redundancy; in AI suggestions, designers should personalize information based on individuals’ domain knowledge and avoid providing information that is commonly known among domain experts.
Provide users with diverse options for the presentation type of AI suggestions
Despite the effectiveness of interventions in cultivating appropriate reliance on AI, research in explainable AI has focused on explanations for binary decision-making tasks; however, AI systems often generate information in a limited presentation format and a lack of user control. The varied amount of time investment could result in a different level of caution when processing information from AI. While these lengthy paragraphs might suit the information-seeking habits of vigilant people, they might not be processed carefully by individuals who tend to be buckpassers. As such, developers should consider providing features that enable users to select their own preferred presentation format based on their own information-seeking habits. For example, if users may adjust the length of text and amount of details (i.e. summary of references or concrete examples) used in AI suggestions, vigilant individuals may choose to obtain more comprehensive information whereas buckpassers could request brief explanations.
6. Limitations and Future Work
We acknowledge several limitations of this study. First, the observed behaviors in the design of this study may not generalize to other domains and tasks. Our study was designed to capture participants’ collaborative decision-making behavior with an AI in a task that required them to accurately identify whether a set of nutrition statements are facts or myths. Choosing nutrition as a domain has the advantage that the study was broadly accessible to diverse participants, but inaccuracies have naturally less severe consequences than tasks in, e.g., the medical domain. We additionally chose to conduct an online study with volunteer participants who were offered personalized performance feedback in exchange for study completion. This usually means participants are intrinsically motivated to do well; however, the consequences of mistaking a nutrition statement as true when it was false (and vice versa) are mild when compared to a real-world situation. While similar effects of decision-making patterns may also occur in other domains and general-purpose tasks, these effects could be influenced by task differences. For tasks that are more complex, these effects may be exacerbated by the complexity of the task (Bansal et al., 2019). For tasks that have less objective ground truth, these effects may still persist yet decision outcomes may change based on individuals’ expertise. Future work is needed to confirm how task complexity and outcome severity shape these effects. We encourage researchers to adopt similar study designs in other settings and professional domains to better understand these dynamics.
Secondly, our study did not fully capture individuals’ various reliance outcomes. Prior studies on AI-assisted decision-making investigated individuals’ appropriate reliance (overreliance or underreliance) on AI (Cao et al., 2024; Vasconcelos et al., 2023). These measures were obtained in a study design that involves two-stage decision-making (e.g., individuals are asked to make decisions first before they are exposed to AI suggestions, and then make final decisions after seeing AI suggestions). Our study adopted a design setting different from two-stage decision-making, because we focused on whether participants seek out AI suggestions and how they perceive their reliance on the AI overall. This means that our study outcome is only partially comparable to prior findings on AI overreliance. Specifically, while we observed that participants who scored high in buckpassing tend to perform worse than those who scored low in buckpassing when both were exposed to incorrect AI decisions, this difference was not significant. Therefore, while participants with high buckpassing tendency were significantly more likely to seek out AI suggestions, we cannot draw firm conclusions about how this tendency influences their overall decision performance. Future work is needed to disentangle whether buckpassers tend to overrely on AI in addition to being more likely to seek AI assistance.
Thirdly, our study design did not fully capture users’ holistic experiences with AI systems, such as ChatGPT, through chatbot interfaces since users did not have a choice to see more AI explanations if they were not satisfied with the output. We chose such a design to ensure AI suggestions are consistent for each participant, yet we hope future studies could explore experiment designs that closely mimic interactions with generative AI systems. Meanwhile, our design of asking users’ to indicate their adoption of AI suggestion by providing an option “Use the AI Decision” had very few engagement click despite the high reliance suggested by other measures. We suspect this occurred due to an unintuitive display of options for users to choose from or a motivation to preserve autonomy in making the final decision. In future work, we will also improve our designs of interface features that are more closely matched with users’ behaviors. It’s also worth noting that our design of the study does not provide the opportunity to search for information from the web. In real-world scenarios, users could potentially navigate other information sources for more information before they make their final decisions. Future work could further integrate this option to examine how users’ behaviors are influenced by their decision-making patterns in a more dynamic setting. Notably, while we did not examine the effect of procrastination in our study, future work that integrates more dynamic interactions shall investigate the behaviors of individuals who tend to procrastinate in their decision-making.
Lastly, our sample primarily consisted of participants who were fluent in English and motivated to participate in our study due to our study topic. In future work, we hope to gather a larger and more diverse sample via by providing the study in other languages and various topics to follow up on the present findings. In line with this, one exciting avenue for future work is to investigate how differences in decision-making patterns across countries and cultures impact people’s reliance on an AI. For instance, Mann et al. (1998) had found that people from more hierarchical cultures commonly score high in buckpassing, which could imply that there are cross-cultural differences in reliance, and potentially in overreliance, on an AI.
7. Conclusion
With the increasing use of conversational AI such as ChatGPT for information-seeking and decision-making, it is essential to understand how people vary in their interactions with AI suggestions in the decision-making process. Through an online study, we asked participants (n=810) to evaluate the factuality of nutrition-related statements with the option to seek AI suggestions (decisions and explanations). We found that people who tend to defer decisions to others (buckpassers) are more likely to seek AI suggestions yet spend less time evaluating these suggestions and reported a higher level of reliance on AI when evaluating nutrition information than those scoring low on buckpassing. In contrast, vigilant decision makers tend to more carefully scrutinize the AI’s information than those scoring low on vigilance. Drawing insight from psychology research on decision-making, our study suggests that individuals’ decision-making patterns implicate not only human-human interactions but also human-AI interactions. In particular, these findings expand the current research horizon of AI-assisted decision-making by underscoring the importance of individual cognitive processes. As more AI-driven systems are being developed and integrated into our everyday lives, these findings shed light on the importance of individual cognitive factors, providing new insights for the future development of AI technologies.
Acknowledgements.
This work is funded by the National Science Foundation under Grant No.2230466.
Appendix A Nutrition Statements
Code Statement Nature Presented AI Decision
S1 Drinking water during meals, contributes to weight gain. Myth Myth
S2 The digestion process begins in the mouth. Fact Fact
S3 Fruit should be eaten before meals. Myth Myth
S4 Egg consumption increases blood cholesterol. Myth Myth
S5 Drinking milk is bad for health. Myth Myth
S6 Eating carbohydrates at night leads to an increase in weight gain. Myth Fact
S7 Fat is important to the human body. Fact Fact
S8 Fruit should be eaten after meals. Myth Myth
S9 Fiber intake is important for normal bowel function. Fact Fact
S10 Gluten-free foods are better for health and should, there-fore, be adopted by all. Myth Myth
S11 Cheese consumption is bad for memory. Myth Myth
S12 Coconut oil is healthier than olive oil. Myth Myth
S13 Lactose-free foods are better for health and should, there-fore, be adopted by all. Myth Myth
S14 Children have different nutritional needs than those for adults. Fact Fact
S15 Fruits and vegetables do not contribute to weight gain. Myth Myth
S16 Normal potatoes are more caloric than sweet potatoes. Myth Fact
S17 Diet should be adapted to a person’s blood group. Myth Fact
S18 Not having a balanced and varied diet can lead to the development of multiple diseases. Fact Myth
S19 The alkaline diet allows balancing the acidity in the blood. Myth Myth
S20 Drinking, while fasting, a glass of water with lemon helps in weight loss. Myth Fact
S21 Inadequate eating habits are the third risk factor for the loss of years of healthy life. Fact Fact
S22 Ingesting high amounts of protein helps in the faster formation of muscles. Myth Fact
S23 Pregnant women should be eating for two. Myth Fact
S24 Cold water should not be drunk. Myth Myth
S25 The day should always start with breakfast. Fact Myth
S26 Water is essential to the normal function of all organs. Fact Fact
S27 Soy milk is healthier than cow’s milk. Myth Myth
S28 Orange should not be eaten at the same time as milk or yogurt. Myth Myth
S29 Dairy products should be consumed in between two and three portions per day. Fact Myth
S30 All food additives (E’s) are harmful to health. Myth Myth
Table 7. Nutrition statements and presented AI’s decisions used in the study’s experiment for participants. These statements are adopted from Florença et al. (2021).
This image is a table that displays nutrition statements and the corresponding AI’s decision on whether the statement is a myth or a fact. The table has three columns: Code, Statement, and Nature/Presented AI Decision.The statements cover a range of topics related to nutrition, such as the effects of consuming certain foods or beverages, dietary recommendations, and general health claims. The ”Nature” column indicates whether the statement is a myth or a fact, and the ”Presented AI Decision” column shows whether the AI deemed the statement to be a myth or a fact.The table covers 30 different nutrition statements, with the AI classifying 12 as facts and 18 as myths. Some examples of the statements include ”Drinking water during meals, contributes to weight gain” (classified as a myth), ”The digestion process begins in the mouth” (classified as a fact), and ”Gluten-free foods are better for health and should, there-fore, be adopted by all” (classified as a myth).Overall, this table provides a comprehensive overview of the AI’s assessment of various nutrition-related claims, which could be useful for understanding the system’s knowledge and decision-making process in this domain.
Appendix B The Melbourne Decision-Making Questionnaire
Table 8. The Melbourne Decision-Making Questionnaire by Mann et al. (1997), which we used in our study to assess participants’ decision-making patterns. All items could be answered on a 3-point scale: “Not true for me”, “Sometimes true for me”, or “True for me”.
Vigilance Item
1 I like to consider all of the alternatives.
2 I try to find out the disadvantages of all alternatives.
3 I consider how best to carry out a decision.
4 When making decisions I like to collect a lot of information.
5 I try to be clear about my objectives before choosing.
6 I take a lot of care before choosing.
Buckpassing
7 I avoid making decisions.
8 I do not make decisions unless I really have to.
9 I prefer to leave decisions to others.
10 I do not like to take responsibility for making decisions.
11 If a decision can be made by me or another person I let the other person make it.
12 I prefer that people who are better informed decide for me.
Hypervigilance
13 Whenever I face a difficult decision I feel pessimistic about finding a good solution.
14 I feel as if I am under tremendous time pressure when making decisions.
15 The possibility that some small thing might go wrong causes me to swing abruptly in my preference.
16 I cannot think straight if I have to make a decision in a hurry.
17 After a decision is made I spend a lot of time convincing myself it was correct.
The Melbourne Decision-Making Questionnaire by Mann et al. [44], used to assess participants’ decision-making patterns. Items are answered on a 3-point scale: ’Not true for me,’ ’Sometimes true for me,’ or ’True for me.’ - Vigilance (Items 1–6): Statements focus on considering alternatives, analyzing disadvantages, collecting information, and being deliberate and clear in decision-making. - Buckpassing (Items 7–12): Statements highlight avoiding or deferring decisions, preferring others to decide, and avoiding responsibility. - Hypervigilance (Items 13–17): Statements express feeling pressured, pessimistic, or anxious about decisions, and struggling with rushed or uncertain choices.
B.1. Study Design
To ensure that participants were aware of any inaccurate information we presented to them during the study, we included at the end of the study a page about their performance in the nutrition evaluation task. Specifically, we included statements where they submitted inaccurate answers and the correct answers for them as well as statements where AI suggestions were not accurate.
Figure 5. Example of result page in the study presented to participants: we included an example of our result page regarding how we debriefed participants about their performance in the nutrition evaluation task. Specifically, we included the statements where they submitted inaccurate answers and statements where AI suggestions were inaccurate.
An image showing an example of a result page presented to participants in a study. The page provides feedback on participants’ performance in a food knowledge task, showing that they answered 5 out of 10 questions correctly. It lists the questions they got wrong and the correct answers, such as ”Diet should be adapted to a person’s blood group. The answer is: This is a myth,” and ”The day should always start with breakfast. The answer is: This is a fact.” Additionally, the page explains that ChatGPT was used to generate AI-based explanations and decisions, noting that some may be inaccurate. Specific questions where AI suggestions were inaccurate are highlighted, including topics like dietary adaptation to blood groups and the caloric comparison between normal and sweet potatoes.
B.2. Factor Analysis for Decision-making Patterns
Figure 6. Factor loading diagram confirming the three decision-making patterns based on the Melbourne Decision Making Questionnaire. Items associated with buckpassing load strongly on PA1, items related to vigilance load on PA2, and items linked to hypervigilance load on PA3, supporting the theoretical structure of the questionnaire. The inter-factor correlation between PA1 and PA3 is shown with a value of 0.7. Loadings represent the strength of the relationship between items and factors, with higher values indicating stronger associations.
A factor loading diagram illustrating the three decision-making patterns based on the Melbourne Decision-Making Questionnaire. The diagram includes three latent factors: PA1 (Buckpassing), PA2 (Vigilance), and PA3 (Hypervigilance). Items related to buckpassing (e.g., questions-buckpassing-p1 through p6) have strong loadings on PA1, items related to vigilance (e.g., questions-vigilance-p1 through p6) load on PA2, and items related to hypervigilance (e.g., questions-hypervigilance-p1 through p5) load on PA3. The inter-factor correlation between PA1 and PA3 is depicted with a value of 0.7. Each loading value (ranging from 0.5 to 0.8) represents the strength of the relationship between individual items and their respective factors, with higher values indicating stronger associations.
Appendix C Demographics Information of Participants
Figure 7. Overall distribution of total time spent in the study: we selected time thresholds (240 seconds and 1200 seconds) where the number of participants significantly dropped and then filtered out participants based on these time threshold.
Figure 8. Age distribution of 810 participants.
A histogram showing the age distribution of 810 participants. The x-axis represents age, ranging from 12 to 90 years, while the y-axis represents the count of participants. The distribution is skewed towards younger ages, with the highest count in the 15–25 age range. The number of participants decreases steadily with age, showing a sharp decline after 30 and tapering off significantly beyond 60. Only a few participants are above 70 years old.
Table 9. Sample sizes as well as averages and standard deviations for each decision-making measure across different demographics. The maximum for vigilance and buckpassing is 12, whereas the maximum for hypervigilance is 10.
N Vigilance Hypervigilance Buckpassing
Overall
   Overall 810 9.41 $\pm$ 2.31 4.25 $\pm$ 2.61 4.56 $\pm$ 3.03
Education
   ¡=High School Education 231 9.39 $\pm$ 2.32 4.94 $\pm$ 2.55 5.25 $\pm$ 3.12
   College Education 403 9.34 $\pm$ 2.34 4.12 $\pm$ 2.64 4.49 $\pm$ 3
   Graduate Education 176 9.59 $\pm$ 2.2 3.64 $\pm$ 2.45 3.82 $\pm$ 2.78
Gender
    Female 409 9.43 $\pm$ 2.31 4.52 $\pm$ 2.63 4.69 $\pm$ 3.15
    Male 363 9.36 $\pm$ 2.32 3.83 $\pm$ 2.47 4.33 $\pm$ 2.84
    Other 38 9.66 $\pm$ 2.13 5.5 $\pm$ 2.94 5.47 $\pm$ 3.27
Age Group
   Under 18 155 9.30 $\pm$ 2.35 5.32 $\pm$ 2.43 5.59 $\pm$ 3.09
   18–24 215 9.69 $\pm$ 2.22 4.64 $\pm$ 2.52 5.01 $\pm$ 2.97
    25–34 174 9.32 $\pm$ 2.51 4.37 $\pm$ 2.6 4.87 $\pm$ 3.03
    35–44 112 9.53 $\pm$ 2.24 3.49 $\pm$ 2.48 3.7 $\pm$ 2.78
    45–54 88 9.03 $\pm$ 2.19 3.05 $\pm$ 2.45 3.03 $\pm$ 2.58
    Over 55 66 9.26 $\pm$ 2.14 3.14 $\pm$ 2.42 3.39 $\pm$ 2.6
Appendix D Correlation Matrix for All Variables
Table 10. Correlation matrix of demographic variables. *** significance at the .001 level; ** significance at the .01 level; * significance at the .05 level; . significance at the 0.1 level.
B V H A HE CE GE DKE
Buckpassing (B) 1
Vigilance (V) -0.05 1
Hypervigilance (H) 0.59*** 0.05 1
Age (A) -0.27*** -0.05 -0.28*** 1
Less than high school education (HE) 0.14*** -0.00 0.17*** -0.35*** 1
Pursuing or Have obtained college education (CE) -0.02 -0.03 -0.05 0.11** -0.63*** 1
Pursuing or Have obtained graduate education (GE) -0.13*** 0.04 -0.12*** 0.25*** -0.33*** -0.52*** 1
Domain Knowledge (Expert) (DKE) -0.20*** 0.07* -0.20*** 0.22*** -0.15*** 0.00 0.16*** 1
Figure 9. Distribution of Conditional Agreement with AI Decision Rate: conditional agreement rate is participants’ agreement rate conditioned on whether they chose to see AI decisions. There are 395 out of 810 participants (49%) who agreed with AI decisions when they chose to see them.
Figure 10. Overall performance breakdown among participants with different levels of buckpassing: participants who chose to see AI decisions showed higher accuracy when the AI’s decisions were correct and lower accuracy when the AI’s decisions were incorrect, across all levels of buckpassing. When exposed to incorrect AI decisions, participants with high buckpassing scores (45%) showed lower accuracy than those with low buckpassing scores (49%). When exposed to correct AI decisions, participants with high buckpassing scores (84%) performed similarly to those with low buckpassing scores (85%). The n indicates the number of samples in each group and condition.
Appendix E Additional Regression Analyses
Table 11. Mixed-effects logistic regression results predicting See_AI_Explanations: no statistically significant effect of decision-making patterns is observed.
Fixed Effects Estimate Std. Error z value p-value
(Intercept) -1.9223 0.32228 -5.965 << 0.001 ***
buckpassing 0.48502 0.31146 1.557 0.119
vigilance -0.02133 0.32836 -0.065 0.948
hypervigilance -0.18182 0.29880 -0.608 0.543
Table 12. Mixed-effects logistic regression result predicting whether participants correctly evaluate whether each nutrition statement is a fact or a myth: result indicates no effects of decision-making patterns. However, participants who chose to see AI decisions and see AI explanations were more likely to have incorrect responses.
Predictors Odds Ratios CI
(Intercept) 5.56∗∗∗ 2.86 – 10.79
Buckpassing 1.00 0.98 – 1.03
Hypervigilance 0.98 0.95 – 1.01
Vigilance 0.99 0.96 – 1.02
Age 1.00 0.99 – 1.00
Domain Knowledge (Average) 0.88 0.77 – 1.02
Domain Knowledge (Below Average) 0.80 0.65 – 0.97
Pursuing or Have obtained college education 1.10 0.94 – 1.29
Pursuing or Have obtained graduate education 1.20 0.99 – 1.46
N Participant 810
N statement 30
Observations 8100
Marginal R2 / Conditional R2 0.003 / 0.432